Please help me out in this sequence of finding r! Thank you so much!
Find r given a(sub 0) is -20. and a(sub 4) = -2500
Thank you for your time in helping me!

Answer by stanbon(75887)

Find r given a(sub 0) is -20. and a(sub 4) = -2500
----
General Form: a(n) = a0*r^(n-1)
----
a(4) = (-20)*r^(3) = -2500
r^3 = 125
r = 5
===================================
Cheers,
Stan H.